# Unleashing Expert Opinion from Social Media for Stock Prediction

### v1

@@ 2. Each extracted subsequence is processed by its own LSTM network:
3. The final representation is obtained by first extracting the last hidden state from each scale ($\mathbf{H}^{(i)}_{t=L/s_{i}}$), averaging these states across all scales, applying layer normalization, and then transforming through an MLP to predict return ratios:
@@
This multi-scale architecture allows the model to capture both short-term and long-term dependencies in the input sequence and produces return ratio estimates for each stock.
@@ V-B2 Loss Function
where $\hat{r}_{i}$ and $r_{i}$ are the predicted and actual return ratios for the $i$-th sample, and $\overline{\hat{r}}$ and $\overline{r}$ are the mean predicted and actual return ratios, respectively. The IC loss function $\mathcal{L}_{IC}$ is defined as:
@@ V-C Dual Graph Attention Network
For each stock on each day, the generated return ratio which from the pre-trained model offers a relative measure of expected returns is combined with two key features:
1. Pseudo label availability: A binary feature indicating whether an expert prediction exists for that day (1 if an expert prediction exists, 0 otherwise).
2. Expert signal: When an expert signal is available, the previously calculated return ratio prior is used. If no expert prediction is made for a stock on a given day, its prior is set to 0.
These three features are combined to form a comprehensive feature set for each stock-day pair. The pseudo label availability helps the model understand when to rely more heavily on the pre-trained output, while the return ratio prior, when available, acts as a corrective signal to refine the baseline prediction. Utilizing this feature set, we propose using the graph attention network to propagate the sparse expert-derived signals and refine the initial estimates from the temporal pre-training model.
For the construction of the graph, we create two types of graphs to represent stock relationships: an industry graph and a dynamically updated correlation matrix graph. For the industry graph, each node represents a stock, and edges are established between nodes if the corresponding stocks belong to the same Global Industry Classification Standard (GICS) sector. The correlation matrix graph is more dynamic, reflecting changes in the stock relationships based on the previous 30 days of price movements. The following is the process for constructing the dynamic correlation matrix graph:
1. Data Preparation: Collect the past 30 days of stock price data.
2. Correlation Calculation: We calculate the correlation coefficients for pairs of stocks and compute the correlation matrix based on the past 30-day window.
3. Graph Construction: Identify positively correlated stock pairs. If the correlation coefficient between two stocks is greater than the set threshold ($\theta_{1}=0.77$ as the general threshold and $\theta_{2}=0.67$ as the threshold for stocks with expert-derived pseudo labels), an edge will be created between this stock pair in the graph. Notably, for stocks associated with expert-derived pseudo labels, we apply a lower correlation threshold, allowing these labeled stocks to influence their connected peers more strongly. The leftmost diagram in Figure 3 illustrates the construction process of both correlation and industry graphs. In the correlation graph example, an edge exists between nodes 2 and 5 as their correlation coefficient (0.68) exceeds $\theta_{2}$, with node 2 containing expert-derived pseudo labels, also known as expert signals. However, despite having a higher correlation coefficient than nodes 2 and 5, nodes 3 and 5 are not connected because their correlation coefficient fails to reach $\theta_{1}$, as neither node contains expert signals.
@@
    Graph Attention Layers: For each graph (industry and correlation), we employ Graph Attention Networks (GAT) to learn node representations. At each layer, a node $v$ aggregates information using attention mechanisms from its neighborhood $\mathcal{N}(v)$, where $\mathcal{N}(v)$ includes both the adjacent nodes and node $v$ itself (i.e., $v\in\mathcal{N}(v)$). The attention coefficients $\alpha_{vu}$ between node $v$ and each node $u\in\mathcal{N}(v)$ are computed as:
@@
    where $q_{ind},q_{cor}\in\mathbb{R}^{d_{hidden}}$ are learnable attention vectors. The attention weights are normalized using softmax:
    $[\beta_{ind},\beta_{cor}]=\text{softmax}([\alpha_{ind},\alpha_{cor}])\in\mathbb{R}^{2\times N}$ (11)
    The fused features are then computed as:
@@
    where $\beta_{ind}^{\top},\beta_{cor}^{\top}\in\mathbb{R}^{N\times 1}$ are the transposed attention weights, $\otimes$ denotes the Kronecker product with $\mathbf{1}_{d_{hidden}}\in\mathbb{R}^{d_{hidden}}$ (broadcasting to $\mathbb{R}^{N\times d_{hidden}}$), $\odot$ represents element-wise multiplication, and $H_{fused}$ represents the fused node features that will be used as input for the next graph attention layer or final prediction.
@@
Our dual-graph attention mechanism enables effective information propagation through both industry relationships and correlation-based connections. The model adaptively learns the relative importance of each graph structure for different nodes during the message-passing process, leading to optimal feature fusion at each hop. Leveraging this adaptive fusion mechanism, the combination of dual-graph attention and expert-derived priors allows the model to refine and correct initial predictions. By incorporating these sparse yet valuable expert signals, the synergy between the temporal model’s output and the expert signals enhances the accuracy and robustness of information propagation across all stocks, ultimately improving predictive performance.
@@ VI-A1 Dataset
We accessed this data through the StockTwits API [^2] https://api.stocktwits.com/api/2/streams/symbol/stock.json, specifically using the endpoint to fetch publicly available data related to specific stocks. During collection, we implemented an automatic pagination mechanism in our data retrieval script. This allowed us to capture both the most recent and complete historical data from the forum.
To optimize the efficiency of our data retrieval process, we employed a distributed system approach using proxy services, enabling multi-threaded, multi-node data collection. This setup significantly accelerated our data gathering efforts; a single node can collect one day’s data from the forum in about 30 minutes. By employing 30 nodes simultaneously, we significantly accelerated the data collection process, retrieving all daily posts within a minute. This substantial increase in speed ensured that we could quickly gather the latest data, enabling timely mining of expert signals to facilitate subsequent model training.
The collected data contains a vast amount of valuable information, with each tweet corresponding to a user’s opinion about a specific stock. For each tweet, the data includes the content, the stock it refers to (including its associated industry, sector, symbol, and the exchange where it is listed), the creation time of the tweet, the user ID, the number of likes it received, the sentiment expressed by the user (either bullish or bearish), and the retweet information. Our data collection spanned from 2010 to the end of 2023, and we have established a database to store all the collected data. This allows users to access various StockTwits data directly from the database without the need to write complex scripts for data extraction. Stock price data was sourced from Polygon.io [^3] https://polygon.io, and fundamental data from EODHD [^4] https://eodhd.com.
For the pre-training of our temporal model, we utilized the collected stock prices and fundamental data spanning from 2010 to 2018. To help enrich our expert signal collection and avoid isolated nodes in the industry graph, we augmented our dataset with 10 highly discussed stocks beyond the index constituents. Our DualGAT model was trained and validated using data from 2019 to 2021 and conducted extensive testing of our DualGAT model on three distinct datasets: the Nasdaq 100 index, S&P 500 index, and the updated StockNet dataset by [61], using data from 2022 to 2023 as the test sets.
@@ VI-A2 Baseline
1. Expert signal baseline: As illustrated in Figure 1, the expert selection methods proposed by Wang et al.[20] in the MFN model, the smart user classification model proposed by Coyne et al. [19], the user centric method proposed by Bouadjenek et al. [27] and the Wining model proposed by Liao et al. [28] were selected as manually designed baseline models to assess the accuracy of the expert signals.
2. Simple message propagation baseline: To demonstrate the effectiveness of the message-passing capability in our designed DualGAT, we employ the simplest form of message passing without graph neural network architectures, denoted as $Wx$, where $W$ is the normalized adjacency matrix derived from either our industry or correlation matrix, and $x$ represents the stock’s return ratio prior mentioned before. Similar to DualGAT, this message-passing method also performs a two-hop propagation. This process helps to diffuse the return ratio prior from nodes (stocks) with expert signals (expert-derived pseudo labels) to adjacent nodes, thereby leveraging the structural connections of the graph to enhance the predictive ability. This baseline is crucial for understanding the ability of the expert signal independently as a trading signal. At the same time, it also serves as the baseline for our DualGAT model that combines traditional financial features.
@@ VI-A3 Metrics
  • ACC (Accuracy): This metric reflects the percentage of correct predictions our model makes regarding the upward or downward trend of stock prices.
  • IC (Information Coefficient) and ICIR: IC measures the correlation between the predicted values and the actual values. It quantifies the ability of the model to make correct relative predictions rather than absolute stock price predictions. ICIR, which adjusts the IC to account for signal volatility, provides a refined measure of predictive performance over time.
  • RIC (Rank Information Coefficient): RIC measures the correlation between the predicted rankings of stocks and their actual rankings. It quantifies the model’s ability to correctly rank stocks relative to each other.
  • AR (Annualized Return): For returns, our strategy involves taking long positions in the top 10% of stocks predicted to have the highest return ratio, and short positions in the 10% predicted to decrease the most. This method is designed to capitalize on the extremes of the market’s movements, thereby maximizing potential returns from market volatility.
  • SR (Sharpe Ratio): The Sharpe Ratio is calculated by dividing the average return earned in excess of the risk-free rate by the standard deviation of return. It provides an adjusted measure of return that accounts for the risk taken, offering insights into the risk-adjusted performance of our investment strategy.
These metrics collectively provide a comprehensive framework for assessing the effectiveness and reliability of our predictive model in a dynamic financial environment.
@@ VI-B Main Results
For the mining of expert signals, Figure 1 illustrates the performance of several expert mining methods across three prediction horizons: T+1, T+3, and T+7. The model labeled “Ours” significantly surpasses other models in accuracy at each time point. At T+1, “Ours” shows an accuracy of 77.26%, which is 18% and 21% higher than the second and third place models (User Centric [27] and SUC [19]). At T+3 and T+7, the accuracy of “Ours” remains better than the other models. These results demonstrate robust short-term predictive strength of our expert tracing system, Although there is a gradual decrease in accuracy as the prediction horizon extends, our model still maintains a significant lead above the baseline (Random Guess) across all evaluated horizons.
@@
To demonstrate the effectiveness of the message-passing capability in our designed model, we compare our DualGAT with the simple message passing method denoted as $Wx$ mentioned above. For the propagation matrix $W$,we utilized the correlation matrix ($W_{cor}$), the industry matrix ($W_{ind}$), and a combination of both. The result in Table I shows that DualGAT outperforms all messgae passing methods on all metrics, indicating DualGAT’s architectural advancements.
@@
Moreover, we demonstrate that DualGAT effectively addresses the sparsity of expert signals. Since the graph attention layer in our neural network corresponds to one hop of propagation, its signal propagation capability is comparable to simple message passing methods (similar to BFS, where edges in the two graphs can spread expert signals). Through empirical analysis using simple message passing methods, we show that a two-hop propagation setting achieves sufficient node coverage, reaching 89.1% of all nodes. DualGAT enhances this propagation process by incorporating learnable attention coefficients, which dynamically adjust the importance of different nodes. Furthermore, attention mechanisms are applied between the industry and correlation graphs, enabling the model to adaptively weigh the contribution of each graph structure during the propagation process, thereby improving the prediction performance.
@@
When comparing our system with other temporal or spatial-temporal models used in time series forcating or stock prediction using the same financial features (price-volume data and fundamental data) on the same test sets shown in Table II, we find that our model not only leads but also significantly outperforms the traditional and more recent models in time series forecasting and stock prediction across all metrics on all datasets. Notably, even for stocks without expert signals, where predictions rely solely on propagated information through our DualGAT architecture, the model still significantly outperforms the baseline models, demonstrating the effectiveness of our signal propagation mechanism. In addition, the results of the pre-trained temporal model MS-LSTM can also be seen in Table II. Since both our model and the baseline model use the same traditional financial features, the significant improvement of our model with the addition of expert signals demonstrates that our expert signal captures valuable information beyond what traditional financial features provide. Overall, these results affirm the superiority of our approach, especially in utilizing expert-derived signals enhanced by graph-based learning.
@@
Mixer [62] 49.81 1.00 1.27 5.88 7.38 0.38 50.72 3.70 2.02 19.64 19.76 1.79 51.06 1.44 1.64 5.14 5.86 0.15
PatchTST [63] 49.58 1.90 1.04 7.22 12.38 1.24 50.36 1.76 1.89 15.19 9.49 0.91 49.85 2.49 2.51 11.35 17.36 1.47
Crossformer [64] 49.34 0.76 0.56 3.14 9.36 0.43 49.28 2.07 0.38 10.04 7.11 0.40 50.95 0.01 -0.05 0.04 -2.11 -0.52
TimesNet [65] 49.74 0.92 0.91 5.13 9.72 0.70 50.73 2.06 1.62 11.53 10.42 0.87 49.03 0.55 0.70 2.19 3.38 -0.06
@@
TimeMixer [66] 50.32 0.71 0.91 4.09 7.68 0.30 50.37 1.40 1.28 10.32 10.78 0.78 50.33 0.76 1.14 4.25 -0.09 -0.41
LSTM-RGCN [67] 49.46 1.10 1.07 6.30 17.73 1.20 49.30 1.59 0.95 15.36 14.56 1.54 49.02 0.90 1.42 4.93 5.24 0.11
SFM [68] 49.24 2.81 2.51 18.37 2.41 2.29 49.30 3.37 3.47 19.85 21.17 2.10 49.01 1.40 1.51 7.38 6.71 0.95
HATR-I [53] 49.38 0.41 0.68 1.90 4.43 0.01 / / / / / / 49.08 3.89 2.67 16.81 21.34 1.64
THGNN [69] 50.75 0.76 0.63 3.73 9.78 0.36 50.79 4.20 4.25 15.36 21.83 1.81 49.59 0.67 0.72 2.76 4.97 0.08
SAMBA [51] 50.68 1.04 0.84 5.59 13.13 0.73 50.58 2.55 1.50 13.84 20.09 1.66 50.30 0.70 0.59 2.55 11.43 0.50
Ours w/o expert signal 57.66 7.83 9.49 35.60 32.55 2.64 54.94 7.15 7.80 36.68 26.93 2.61 53.82 6.29 6.71 26.04 20.41 1.26
Ours w/ expert signal 57.83 7.98 9.69 36.66 32.91 2.69 55.02 8.26 8.29 43.17 29.56 2.92 54.81 7.62 7.78 30.69 28.46 1.87

### v2

@@ 2. Each extracted subsequence is processed by its own LSTM network:
[…]
@@
This multi-scale architecture allows the model to capture both short-term and long-term dependencies in the input sequence and produces return ratio estimates for each stock. The superiority of MS-LSTM can be seen in Supplementary Material S8.
@@ V-B2 Loss Function
where $\hat{r}_{i}$ and $r_{i}$ are the predicted and actual return for the $i$-th sample, and $\overline{\hat{r}}$ and $\overline{r}$ are the mean predicted and actual return, respectively. The IC loss function $\mathcal{L}_{IC}$ is defined as:
@@ V-C Dual Graph Attention Network
For each stock on each day, the return ratio estimate generated from the pre-trained model is combined with two key features:
1. Expert availability: A binary feature indicating whether expert predictions exist for that day (1 if exist, 0 otherwise).
2. Expert signals: When an expert signal is available, the previously calculated expert signal value is used. If no expert prediction is made for a stock on a given day, it is set to 0.
These three features are combined to form a comprehensive feature set for each stock-day pair. Expert availability helps the model understand when to rely more heavily on the pre-trained output, while expert signals, when available, act as corrective signals to refine baseline predictions. Using this feature set, we propose using graph attention networks to propagate sparse expert signals and refine initial estimates from the temporal pre-training model.
For graph construction, we create two types of graphs to represent stock relationships: an industry graph and a dynamically updated correlation matrix graph. For the industry graph, nodes represent stocks, and edges are established between nodes if corresponding stocks belong to the same Global Industry Classification Standard (GICS) sector. The correlation matrix graph is more dynamic, reflecting changes in stock relationships based on the previous 30 days of price movements. The following is the process for constructing the dynamic correlation matrix graph:
1. Data Preparation: Collect the past 30 days of stock price.
2. Correlation Calculation: Calculate the correlation coefficients for pairs of stocks and compute the correlation matrix based on the past 30-day window.
3. Graph Construction: Identify positively correlated stock pairs. If the correlation coefficient between two stocks exceeds the set threshold ($\theta_{1}=0.77$ as general threshold and $\theta_{2}=0.67$ for stocks with expert signals), an edge is created between this stock pair. For stocks with expert signals, we apply a lower correlation threshold, allowing these stocks to influence connected peers more strongly. The rationale for these threshold choices and sensitivity analysis can be found in Supplementary Material S3. The leftmost diagram in Figure 3 illustrates the construction process of both correlation and industry graphs. In the correlation graph example, an edge exists between nodes 2 and 5 as their correlation coefficient (0.68) exceeds $\theta_{2}$, with node 2 containing expert signals. However, despite having a higher correlation coefficient than nodes 2 and 5, nodes 3 and 5 are not connected because their correlation coefficient fails to reach $\theta_{1}$, as neither node contains expert signals.
@@
    Graph Attention Layers: For each graph (industry and correlation), we employ Graph Attention Networks (GAT) to learn node representations. These GAT layers correspond to the pink blocks shown in Figure 3. At each layer, a node $v$ aggregates information using attention mechanisms from its neighborhood $\mathcal{N}(v)$, where $\mathcal{N}(v)$ includes both the adjacent nodes and node $v$ itself (i.e., $v\in\mathcal{N}(v)$). The attention coefficients $\alpha_{vu}$ between node $v$ and each node $u\in\mathcal{N}(v)$ are computed as:
@@
    where $q_{ind},q_{cor}\in\mathbb{R}^{d_{hidden}}$ are learnable attention vectors. This computation corresponds to the connecting lines from the graph nodes to the purple ”Dual Graph Attention” blocks in Figure 3. The attention weights are normalized using softmax:
    $\beta_{ind},\beta_{cor}]=\text{softmax}([\alpha_{ind},\alpha_{cor}])\in\mathbb{R}^{2\times N}$ (11)
    These $\beta$ values correspond to the blue attention coefficients shown in Figure 3. The fused features are then computed as:
@@
    where $\beta_{ind}^{\top},\beta_{cor}^{\top}\in\mathbb{R}^{N\times 1}$ are the transposed attention weights, $\otimes$ denotes the Kronecker product with $\mathbf{1}_{d_{hidden}}\in\mathbb{R}^{d_{hidden}}$ (broadcasting to $\mathbb{R}^{N\times d_{hidden}}$), $\odot$ represents element-wise multiplication, and $H_{fused}$ represents the fused node features that will be used as input for the next graph attention layer or final prediction. This fusion mechanism effectively combines information from both graph structures as visualized by the merging arrows in Figure 3.
@@
Our dual-graph attention mechanism enables effective information propagation through both industry relationships and correlation-based connections. The model adaptively learns the relative importance of each graph structure for different nodes during message-passing, leading to optimal feature fusion at each hop. Leveraging this adaptive fusion mechanism, the combination of dual-graph attention and expert signals allows the model to refine and correct initial predictions. By incorporating these sparse yet valuable signals, the synergy between temporal model output and expert information enhances the accuracy and robustness of information propagation across all stocks, ultimately improving predictive performance.
@@ VI-A1 Dataset
We accessed data through the StockTwits API, using the endpoint to fetch publicly available data related to specific stocks. During collection, we implemented an automatic pagination mechanism in our data retrieval script. This allowed us to capture complete historical data from the forum.
To optimize data retrieval efficiency, we employed a distributed system using proxy services, enabling multi-threaded, multi-node data collection. A single node collects one day’s data in about 30 minutes. Using 30 nodes simultaneously can accelerate collection, retrieving daily posts within a minute. This speed increase ensured quick data gathering, enabling timely expert signal mining for subsequent model training.
The collected data contains valuable information, with each tweet representing a user’s opinion about a specific stock. For each tweet, the data includes content, the referenced stock (industry, sector, symbol, and exchange), creation time, user ID, likes received, sentiment (bullish or bearish), and retweet information. Our data collection spanned from 2010 to 2023, and we established a database storing all collected data. This allows users to access StockTwits data directly without writing complex extraction scripts. Stock price data was sourced from Polygon.io, and fundamental data from EODHD.
For the pre-training of our temporal model, we utilized the collected stock prices and fundamental data spanning from 2010 to 2018. To help enrich our expert signals collection and avoid isolated nodes in the industry graph, we augmented our dataset with the 10 stocks that possessed the most expert signals during the training phase. Our DualGAT model was trained and validated using data from 2019 to 2021 and conducted extensive testing of our DualGAT model on three distinct datasets: the Nasdaq 100 index, S&P 500 index, and the updated StockNet dataset by [www2024learning], using data from 2022 to 2023 as the test sets.
@@ VI-A2 Baseline
1. Expert signals baseline: As illustrated in Figure 1, the expert selection methods proposed by Wang et al.[aaai_expert_based] in the MFN model, the smart user classification model proposed by Coyne et al. [Forecating_stock_prices], the user centric method proposed by Bouadjenek et al. [A_user_centric] and the Wining model proposed by Liao et al. [liao2014winning] were selected as manually designed baseline models to assess the accuracy of the expert signals.
2. Simple message propagation baseline: To demonstrate the effectiveness of the message-passing capability in our designed DualGAT, we employ the simplest form of message passing without graph neural network architectures, denoted as $Wx$, where $W$ is the normalized adjacency matrix derived from either our industry or correlation matrix, and $x$ represents the stock’s expert signal. Similar to DualGAT, this message-passing method also performs a two-hop propagation. This process helps to diffuse the expert signals to adjacent nodes, thereby leveraging the structural connections of the graph to enhance the predictive ability. This baseline is crucial for understanding the ability of the expert signals independently as a trading signal. At the same time, it also serves as the baseline for our DualGAT model.
@@ VI-A3 Metrics
  • ACC: This metric reflects the percentage of correct predictions our model makes regarding the upward or downward trend of stock prices.
  • IC and ICIR: IC measures the correlation between predicted and actual values, quantifying the model’s ability to make correct relative predictions rather than absolute stock price predictions. ICIR adjusts IC for signal volatility, providing a refined measure of predictive performance over time.
  • RIC: RIC measures the correlation between the predicted rankings of stocks and their actual rankings. It quantifies the model’s ability to correctly rank stocks relative to each other.
  • AR: For returns, our strategy involves taking long positions in the top 10% of stocks predicted to have the highest return ratio, and short positions in the 10% predicted to decrease the most. We apply a 4 basis points trading cost (0.04%) to account for transaction costs and market impact. This method is designed to maximize potential returns from market volatility.
  • SR: The Sharpe Ratio measures risk-adjusted returns by dividing excess returns by return volatility, offering insights into the risk-adjusted performance of our investment strategy.
The detailed mathematical definitions of these evaluation metrics can be found in Supplementary Material S4. These metrics collectively provide a comprehensive framework for assessing the effectiveness and reliability of our predictive model in a dynamic financial environment.
@@ VI-B Main Results
For the mining of expert signals, Figure 1 illustrates the performance of several expert identification methods across three prediction horizons: T+1, T+3, and T+7. The model labeled “Ours” significantly surpasses other models in accuracy at each time point. At T+1, “Ours” shows an accuracy of 72.8%, which is 13% and 15% higher than the second and third place models (User Centric [A_user_centric] and SUC [Forecating_stock_prices]). At T+3 and T+7, our method also demonstrates substantial leads over existing approaches. To further validate the robustness of our expert identification method, we applied it to different markets and social media platforms. We tested our approach on Xueqiu, which is the largest stock discussion platform in China with a stock pool comprising all stocks from the Shanghai and Shenzhen stock exchanges. Our method also substantially outperforms the leading alternative expert identification method from prior work, achieving approximately 10% higher accuracy on this dataset. Detailed comparisons can be found in Supplementary Material S5. These results demonstrate that our expert identification method is highly effective across different social media sources and diverse market environments. Moreover, the consistent superior performance across various prediction horizons validates the robustness and reliability of our approach for practical investment applications. Additionally, we conducted an experiment to compare the performance of signals derived solely from identified true experts versus those from inverse experts, with the detailed analysis presented in Supplementary Material S6.
@@
To demonstrate the effectiveness of the message-passing capability in our designed model, we compare our DualGAT with the simple message passing method denoted as $Wx$ mentioned above. For the propagation matrix $W$,we utilized the correlation matrix ($W_{cor}$), the industry matrix ($W_{ind}$), and a combination of both. The result in Table I shows that DualGAT outperforms all message passing methods on all metrics, indicating DualGAT’s architectural advancements.
@@
Moreover, we demonstrate that DualGAT effectively addresses the sparsity of expert signals. Since the graph attention layer in our neural network corresponds to one hop of propagation, its signal propagation capability is comparable to simple message passing methods (similar to BFS, where edges in the two graphs can spread expert signals). Through empirical analysis using simple message passing methods, we show that a two-hop propagation setting achieves sufficient node coverage, reaching 89.1% of all nodes. DualGAT enhances this propagation process by incorporating learnable attention coefficients, which dynamically adjust the importance of different nodes. Furthermore, attention mechanisms are applied between the industry and correlation graphs, enabling the model to adaptively weigh the contribution of each graph structure during the propagation process, thereby improving the prediction performance. Our examination of the learned attention weights reveals substantial variation across different stocks between the two graphs, providing compelling evidence of the model’s adaptive behavior. Detailed attention weight analysis can be found in Supplementary Material S7.
@@
When comparing with other temporal or spatial-temporal models used in time series forecasting or stock prediction shown in Table II, we find that our model significantly outperforms the traditional and more recent models in time series forecasting and stock prediction across all metrics on all datasets. Notably, our signal propagation mechanism through our DualGAT architecture proves highly effective. Even when evaluated solely on the subset of stocks without direct expert signals (as shown in the ‘Ours w/o expert signal’ row), it still significantly outperforms all baseline models, demonstrating successful signal propagation from the small fraction of stocks with expert signals to the broader stock universe.
To further validate robustness, we conducted out-of-sample testing across distinct market regimes spanning COVID periods (2019-2023). Our model consistently outperforms across these conditions, showing strong generalization capabilities. Detailed evaluation setup and cumulative return comparisons are in Supplementary Material S8.
Since both our model and the baseline model use the same traditional financial features, the significant improvement of our model with the addition of expert signals demonstrates that our expert signals capture valuable information beyond what traditional financial features provide. Overall, these results affirm the superiority of our approach, especially in utilizing expert signals enhanced by graph-based learning.
@@
Mixer [tolstikhin2021mlp] 49.81 1.00 1.27 5.88 7.38 0.38 50.72 3.70 2.02 19.64 19.76 1.79 51.06 1.44 1.64 5.14 5.86 0.15
PatchTST [nie2022time] 49.58 1.90 1.04 7.22 12.38 1.24 50.36 1.76 1.89 15.19 9.49 0.91 49.85 2.49 2.51 11.35 17.36 1.47
Crossformer [zhang2023crossformer] 49.34 0.76 0.56 3.14 9.36 0.43 49.28 2.07 0.38 10.04 7.11 0.40 50.95 0.01 -0.05 0.04 -2.11 -0.52
TimesNet [wutimesnet] 49.74 0.92 0.91 5.13 9.72 0.70 50.73 2.06 1.62 11.53 10.42 0.87 49.03 0.55 0.70 2.19 3.38 -0.06
@@
TimeMixer [wangtimemixer] 50.32 0.71 0.91 4.09 7.68 0.30 50.37 1.40 1.28 10.32 10.78 0.78 50.33 0.76 1.14 4.25 -0.09 -0.41
LSTM-RGCN [li2021modeling] 49.46 1.10 1.07 6.30 17.73 1.20 49.30 1.59 0.95 15.36 14.56 1.54 49.02 0.90 1.42 4.93 5.24 0.11
SFM [zhang_stock_2017] 49.24 2.81 2.51 18.37 2.41 2.29 49.30 3.37 3.47 19.85 21.17 2.10 49.01 1.40 1.51 7.38 6.71 0.95
HATR-I [wang2022hatr] 49.38 0.41 0.68 1.90 4.43 0.01 / / / / / / 49.08 3.89 2.67 16.81 21.34 1.64
THGNN [xiang2022temporal] 50.75 0.76 0.63 3.73 9.78 0.36 50.79 4.20 4.25 15.36 21.83 1.81 49.59 0.67 0.72 2.76 4.97 0.08
SAMBA [samba] 50.68 1.04 0.84 5.59 13.13 0.73 50.58 2.55 1.50 13.84 20.09 1.66 50.30 0.70 0.59 2.55 11.43 0.50
Ours w/o expert signals 57.66 7.83 9.49 35.60 32.55 2.64 54.94 7.15 7.80 36.68 26.93 2.61 53.82 6.29 6.71 26.04 20.41 1.26
Ours w/ expert signals 57.83 7.98 9.69 36.66 32.91 2.69 55.02 8.26 8.29 43.17 29.56 2.92 54.81 7.62 7.78 30.69 28.46 1.87